### 12 Dec 2023 v1

@@ 3.4.1 Scoring Function
4:Output prediction set: $\sum_{i=1}^{k^{*}}(\hat{\pi}_{(i)}(x+1)+\lambda\textbf{1}[j>k_{\text{reg}}])\geq\hat{q}$
Results presented in Table 2 show that RASP produces an average prediction set size smaller than that generated by APS. While smaller prediction set sizes are typically preferred as they imply greater certainty, this characteristic may pose challenges, particularly in instances of misclassified samples where the regularization term penalizes larger set sizes that show more uncertainty, as shown in Figure 4
@@
ASP 2.84 5.54 4.19
RASP 1.19 2.35 1.77
Table 2: Comparison between the average prediction set size produced by APS and RAPS algorithms for correctly classified samples and misclassified samples.
@@
Figure 4: Comparing the effect of APS Vs. RAPS on uncertainty assignments for both correctly classified samples and misclassified samples.
@@ 3.4.2 Confidence Level
The confidence level ($1-\alpha$) is a user-defined parameter, representing the probability of predicting the true label within the generated prediction set. The value of $\alpha$ is also used to compute the empirical quantile that splits the output scores into conforming and nonconforming scores. According to Equation 8, increasing $\alpha$ reduces the value of the quantile. at $\alpha=1.0$, the value of the quantile is 00, hence the prediction set will produce empty sets of predictions. To further investigate the impact of increasing the value of $\alpha$, we examined the algorithm’s uncertainty assignment through the samples of misclassified data. As illustrated in Figure 5, it is evident that CP assigns high uncertainty values to a higher count of samples at $\alpha=0.1$. This is an expected behavior as the samples are wrongly classified, and hence they belong to the uncertain category. As the value of $\alpha$ increases, and the quantile value decreases, leading to a reduction in the number of samples generated with high uncertainty. Consequently, the count of certain samples increases, and eventually, those samples turn into empty prediction sets at higher $\alpha$ values.
Figure 5: The relationship between the size of the calibration set and the empirical coverage of conformal prediction algorithm
@@ 3.4.3 Calibration set size
Figure 6: The relationship between the size of the calibration set and the empirical coverage of conformal prediction algorithm
@@ 3.5 Uncertainty Quantification with Out-of-Distribution Data
To assess the performance of the three uncertainty quantification methods, Out-of-Distribution (OOD) data is used for testing. OOD is a term that is generally used to refer to the distribution mismatch between the source domain (training data) and the target domain (inferencing data) [30, 31]. Generally, there are two categories of distributional shifts categorized based on the source of discrepancy: semantic shift, where the labels of the target data are different than the source data, and covariate shift, where shifts occur in the input domain, but labels remain unchanged.
In the context of skin lesion classification, we have selected the Dermofit Image Library (DMF) [32] dataset to evaluate the performance of uncertainty quantification algorithms. The DMF dataset contains images from real patients, featuring seven different types of skin lesions. Notably, these lesion types align with those found in the HAM10000 dataset, which was utilized to train the core model. Despite having similar labels for both datasets, each dataset is captured using different medical devices in distinct facilities. This divergence in settings introduces a covariate shift in the distribution of both datasets, presenting a noteworthy challenge for the underlying model. The distribution shift is visualized through a dimensionality reduction t-SNE plot of both datasets, as depicted in Figure 7.
@@
The assigned uncertainty values during inference are recorded to analyze the reaction of each method to shifted data. The uncertainty distributions are shown in Figure 8. Ideally, high uncertainty values should be assigned to samples drawn from distributions different from those used during training. Figure 8 reveals that MCD exhibits inferior performance compared to both EDL and CP. While EDL effectively assigns high uncertainty values to a larger number of samples, CP achieves the best overall performance, where all the samples are assigned a high uncertainty value, indicating that the network is highly uncertain with the predictions.
Figure 8: Histogram distribution of uncertainty estimations by Monte Carlo dropout, evidential deep learning, and conformal prediction algorithms. Dataset with covariate shift is used to assess the performance of the three methods.
@@ 4 Conclusion
In CP, the assumption that the calibration dataset is drawn from an independent and identical distribution (i.i.d) as the training dataset poses a limitation, particularly in real-world scenarios where models encounter diverse data. This assumption becomes restrictive as models are exposed to a broad spectrum of data, making the i.i.d assumption challenging to consistently uphold. While the reported results obtained from the OOD experiment showcase superior CP performance compared to other methods, it’s noteworthy that the testing dataset used doesn’t exhibit significant diversity from the training dataset. This limitation prompts consideration for future research directions, emphasizing the exploration of domain adaptation and generalization techniques to enhance model performance under domain shifts.

### 16 Mar 2024 v2

@@ 3.2.1 Scoring Function
4:Output prediction set: $\sum_{i=1}^{k^{*}}(\hat{\pi}_{(i)}(x_{n+1})+\lambda\textbf{1}[j>k_{\text{reg}}])\leq\hat{q}$
Results presented in Table 3 show that RAPS produces an average prediction set size smaller than that generated by APS. While smaller prediction set sizes are typically preferred as they imply greater certainty, this characteristic may pose challenges, particularly in instances of misclassified samples where the regularization term penalizes larger set sizes that show more uncertainty, as shown in Figure 5
@@
APS 2.84 5.54 4.19
RAPS 1.19 2.35 1.77
Table 3: Comparison between the average prediction set size produced by APS and RAPS algorithms for correctly classified samples and misclassified samples. Results illustrate that RAPS is capable of generating a smaller average set size compared to APS
@@
Figure 5: Comparing the effect of APS Vs. RAPS on uncertainty assignments for both correctly classified samples and misclassified samples. The figure illustrates that smaller prediction set sizes produced by RAPS, often indicating lower uncertainty, are generally preferred. However, challenges arise in cases of misclassified samples where larger set sizes, indicating increased uncertainty, may incur penalties from the regularization term
@@ 3.2.2 Confidence Level
The confidence level ($1-\alpha$) is a user-defined parameter, representing the probability of predicting the true label within the generated prediction set. The value of $\alpha$ is also used to compute the empirical quantile that splits the output scores into conforming and nonconforming scores. According to Equation 8, increasing $\alpha$ reduces the value of the quantile. at $\alpha=1.0$, the value of the quantile is 00, hence the prediction set will produce empty sets of predictions. To further investigate the impact of increasing the value of $\alpha$, we examined the algorithm’s uncertainty assignment through the samples of misclassified data. As illustrated in Figure 6, it is evident that CP assigns high uncertainty values to a higher count of samples at $\alpha=0.1$. This is an expected behavior as the samples are wrongly classified, and hence they belong to the uncertain category. As the value of $\alpha$ increases and the quantile value decreases, leading to a reduction in the number of samples generated with high uncertainty. Consequently, the count of certain samples increases, and eventually, those samples turn into empty prediction sets at higher $\alpha$ values.
Figure 6: The relationship of varying the value of $\alpha$ that corresponds to the confidence level and the count of uncertain samples, certain samples, and empty sets. The figure illustrates that increasing the value of $\alpha$ results in producing more empty sets ($\alpha$ =1).
@@ 3.2.3 Calibration set size
Figure 7: The relationship between the size of the calibration set and the empirical coverage of conformal prediction algorithm. As illustrated in the Figure, It is advisable to incorporate a large number of samples in the calibration data to mitigate fluctuations in coverage and ensure consistent performance.
4 Discussion
In light of the conducted experiments, the obtained results provide valuable insights and analysis into uncertainty prediction techniques. In high-stakes domains like clinical decision-making, where the ramifications of misclassification can be significant, it is imperative to accompany deep learning algorithms’ classification outputs with confidence intervals. Our findings underscore the importance of employing robust uncertainty quantification methods, which yield valuable insights crucial for informed decision-making. Particularly in instances of high uncertainty outputs, involving human expertise in the evaluation process becomes indispensable, augmenting the algorithm’s assessment with human judgment and expertise.
Before discussing the advantages and drawbacks of each technique, it is imperative to emphasize that predictive uncertainty should not be misconstrued as a definitive indicator of classification output. Instead, it serves as supplementary information that enriches the model’s robustness. It’s typical to find correctly classified samples in the dataset with a mix of both high and low uncertainty values. The variability in uncertainty values is a result of the differences in difficulty levels among individual samples. Ideally, the uncertainty value of the wrongly classified samples should be consistently higher, indicating less confidence in the network’s output. As observed in Figure 2, CP demonstrates a notable trend wherein a significant proportion of misclassified samples are attributed high uncertainty values. Conversely, upon analyzing the outcomes produced by MCD and EDL, we note a more uniform distribution of uncertainty values for misclassified samples. This observation hints at a broader spectrum of uncertainty values, indicating a higher degree of variability in the predictive confidence levels generated by these algorithms.
Compared with MCD and EDL, CP stands out for its simplicity in implementation. The algorithm operates seamlessly, eliminating the need for retraining procedures or modifications to the model’s architecture. This crucial feature facilitates the quantification of uncertainties and ensures that the accuracy of the original model remains unaffected. CP, however, requires the calibration dataset to be independent and identical to the training data. This assumption limits the operation of CP in real-world applications, where the tight assumption often does not hold. In MCD, little modifications to the model architecture might be necessary to include dropout layers if not present. Nevertheless, MCD requires running multiple forward passes, contributing to an increased computational load. Moreover, it is crucial to note that high dropout rates have the potential to compromise accuracy, emphasizing the importance of carefully tuning dropout parameters to strike a balance between uncertainty quantification and model performance. Additionally, some studies show that MCD tends to output ill-calibrated uncertainty values that can be unreliable for uncertainty quantification [35]. EDL, on the other hand, generates sharp uncertainty distribution for the correctly classified samples, but a broad distribution with wrongly classified samples, making it less reliable. The implementation of EDL is relatively more complex when compared to CP and MCD. It requires modifying the core model, changing the training loss function, and retraining the modified model from scratch. Table 4 lists a summarization of the advantages and limitations of each of the discussed uncertainty quantification methods.
Table 4: Comparison of the advantages and limitations of the uncertainty quantification methods
UQ Method Advantages Limitations
EDL End-to-end training Architecture adjustments; Sensitivity to hyperparameters
MCD Simple Implementation High Computational Overhead
CP Distribution free; Simple Implementation Exchangability assumption of calibration data
@@ 4.1 Uncertainty Quantification with Out-of-Distribution Data
To assess the performance of the three uncertainty quantification methods, Out-of-Distribution (OOD) data is used for testing. OOD is a term that is generally used to refer to the distribution mismatch between the source domain (training data) and the target domain (inferencing data) [36]. Generally, there are two categories of distributional shifts categorized based on the source of discrepancy: semantic shift, where the labels of the target data are different than the source data, and covariate shift, where shifts occur in the input domain, but labels remain unchanged.
In the context of skin lesion classification, we have selected the DMF [28] dataset to evaluate the performance of uncertainty quantification algorithms. The DMF dataset contains images from real patients, featuring seven different types of skin lesions. Notably, these lesion types align with those found in the HAM10000 dataset, which was utilized to train the core model. Despite having similar labels for both datasets, each dataset is captured using different medical devices in distinct facilities. This divergence in settings introduces a covariate shift in the distribution of both datasets, presenting a noteworthy challenge for the underlying model. The distribution shift is visualized through a dimensionality reduction t-SNE plot of both datasets, as depicted in Figure 8.
@@
The assigned uncertainty values during inference are recorded to analyze the reaction of each method to shifted data. The average uncertainty value generated by each algorithm is reported in Table 5, while the uncertainty distributions are shown in Figure 9. Ideally, high uncertainty values should be assigned to samples drawn from distributions different from those used during training. Figure 9 reveals that MCD exhibits inferior performance compared to both EDL and CP. While EDL effectively assigns high uncertainty values to a larger number of samples, CP achieves the best overall performance, where all the samples are assigned a high uncertainty value, indicating that the network is highly uncertain with the predictions.
UQ algorithm Average Uncertainty
MCD 0.86
EDL 0.51
CP 0.08
Table 5: Average Uncertainty Values for Different uncertainty quantification algorithms. The core network is trained on HAM10000 dataset and evaluated on DMF dataset as an OOD example. Results illustrate that CP achieves the best performance in terms of OOD detection.
Figure 9: Histogram distribution of uncertainty estimations by Monte Carlo dropout, evidential deep learning, and conformal prediction algorithms. A dataset with covariate shift is used to assess the performance of the three methods.
@@ 5 Conclusion
In CP, the assumption that the calibration dataset is drawn from an independent and identical distribution (i.i.d) as the training dataset poses a limitation, particularly in real-world scenarios where models encounter diverse data. This assumption becomes restrictive as models are exposed to a broad spectrum of data, making the i.i.d assumption challenging to consistently uphold. While the reported results obtained from the OOD experiment showcase superior CP performance compared to other methods, it’s noteworthy that the testing dataset used doesn’t exhibit significant diversity from the training dataset. This limitation prompts consideration for future research directions, emphasizing the exploration of domain adaptation and generalization techniques to enhance model performance under domain shifts. In addition, it is worth acknowledging that the presence of an imbalanced dataset can present challenges that reflect real-world situations. Specifically, the effectiveness of uncertainty quantification methods is intricately tied to the performance of the core network. Therefore, it’s essential to recognize that a naturally balanced dataset could potentially enhance the overall performance of these approaches.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
All data and scripts are available at: https://github.com/jfayyad/UQ